OpenStudy (anonymous):

(6 + sqrt 3)(6 - sqrt 3) = ?

OpenStudy (hoblos):

(a+b)(a-b) = a^2 - b^2 use this formula

OpenStudy (experimentx):

6^2 - 3 = 33

OpenStudy (anonymous):

33